OpenStudy (anonymous):

if f(x)= (3x-10)/(9x+9), what is f^-1

hero (hero):

I already posted the solution to that

hero (hero):

hint: y = f(x)

OpenStudy (anonymous):

oh sorry i came up with the same solution and it doesnt workk

OpenStudy (dumbcow):

multiply both sides by (9x+9) --> (9x+9)y = 3x-10 9xy +9y = 3x-10 9xy -3x = -9y-10 x(9y-3) = -9y-10 x = -(9y+10)/(9y-3) = f^-1(x)

OpenStudy (anonymous):

thanks DC

hero (hero):

That's the same thing I posted.

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

help with this one g(x) = 7x^3-3, what is g^-1 (x) im not sure what to do with exponent

hero (hero):

Post it as a separate question.